### v1

Abstract
Background: Global healthcare systems face critical challenges from increasing patient volumes and limited consultation times, with primary care visits averaging under 5 minutes in many countries. Pre-consultation processes—encompassing triage and structured history-taking—offer potential solutions but remain limited by passive interaction paradigms and context management challenges in existing AI systems.
Objective: This study introduces a hierarchical multi-agent framework that transforms passive medical AI systems into proactive inquiry agents through autonomous task orchestration, addressing the limitations of current pre-consultation technologies.
Methods: We developed an eight-agent architecture with centralized control mechanisms for dynamic medical consultation. The framework decomposes pre-consultation into four primary tasks: Triage ($T_{1}$), History of Present Illness collection ($T_{2}$), Past History collection ($T_{3}$), and Chief Complaint generation ($T_{4}$), with $T_{1}$$T_{3}$ further divided into 13 domain-specific subtasks. A Controller agent coordinates specialized agents through dynamic completion assessment, adaptive prompt generation, and hierarchical task management. The system was evaluated on 1,372 validated electronic health records from a Chinese medical platform, with performance assessed across multiple foundation models (GPT-OSS 20B, Qwen3-8B, Phi4-14B).
Results: The framework achieved 87.0% accuracy for primary department triage and 80.5% for secondary department classification. Task completion rates reached 98.2% using agent-driven scheduling versus 93.1% with sequential processing. Clinical quality scores from 18 physicians averaged 4.56 for Chief Complaints, 4.48 for History of Present Illness, and 4.69 for Past History (5-point scale). The system demonstrated robust performance across different language models without task-specific fine-tuning, completing consultations within 12.7 rounds for $T_{2}$ and 16.9 rounds for $T_{3}$.
Conclusions: The hierarchical multi-agent framework successfully enables proactive, structured medical inquiry through intelligent coordination, significantly outperforming conventional passive approaches. The model-agnostic architecture maintains high performance across different foundation models while preserving data privacy through local deployment. These findings demonstrate the potential for autonomous AI systems to enhance pre-consultation efficiency and quality in clinical settings.
keywords:
[…]
1 Introduction
The global healthcare system confronts unprecedented challenges from escalating patient volumes, aging populations, and persistent imbalances between medical demand and workforce capacity. This crisis manifests particularly in consultation durations: systematic reviews across 67 countries reveal that primary care visits last 5 minutes or less in nations representing half the world’s population [10]. In China, consultation times as brief as 4.3 minutes have been documented [21]—durations insufficient for ensuring diagnostic quality and comprehensive patient care. These temporal constraints not only elevate physician burnout risks but fundamentally compromise healthcare delivery quality [11, 16].
In this paper, we focus on pre-consultation, which we define as the process of guiding patients toward appropriate clinical departments (Triage) and systematically collecting essential medical history (History Taking) prior to formal consultation. Pre-consultation has emerged as a critical workflow innovation to alleviate systemic pressures: it enhances diagnostic efficiency, shortens consultation times, and reduces unnecessary clinical visits. However, traditional manual navigation approaches merely redistribute workload from physicians to nursing staff without reducing overall system burden, while also suffering from inconsistent documentation quality and fatigue-induced errors [26, 8].
Recent advances in large language models (LLMs) have enabled the development of intelligent pre-consultation systems with sophisticated dialogue capabilities. Multi-agent frameworks have demonstrated particular promise in healthcare applications, with specialized agents collaborating to enhance diagnostic reasoning [9] and achieving high accuracy in clinical advice provision [1]. However, existing systems exhibit three fundamental limitations: (1) passive interaction paradigms that respond to patient inputs rather than proactively guiding structured inquiry, failing to replicate physician-directed questioning essential for diagnostic quality; (2) context window limitations causing up to 39% performance degradation in extended conversations due to “loss-in-middle-turns” phenomena [12], compromising the 10–20 conversational exchanges required for comprehensive medical evaluation; and (3) absence of hierarchical control mechanisms preventing balanced inquiry management across multiple medical domains, resulting in systems that either over-focus on details or maintain rigid questioning sequences.
To address these challenges, we propose a novel hierarchical multi-agent framework that transforms passive medical AI systems into proactive inquiry agents through autonomous task orchestration and intelligent coordination. Our approach introduces a central Controller that coordinates specialized agents to achieve three core innovations: (1) dynamic subtask completion assessment that monitors information gathering progress across 13 predefined medical domains using medically-informed evaluation criteria, (2) adaptive prompt generation that synthesizes patient responses with accumulated consultation context to formulate clinically optimal follow-up questions, and (3) hierarchical task management that balances macroscopic diagnostic progression with microscopic symptom detail collection through priority-based orchestration.
The framework formalizes the pre-consultation system into four primary tasks: Triage ($T_{1}$), History of Present Illness collection ($T_{2}$), Past History collection ($T_{3}$), and Chief Complaint generation ($T_{4}$). Among them, $T_{1}$, $T_{2}$, and $T_{3}$ are further decomposed into several domain-specific subtasks (see the table below) to support structured and comprehensive information acquisition. Through the collaboration of multiple agents—including the Monitor, Inquirer, and Controller—the system establishes an intelligent navigation framework that integrates authentic clinical logic with autonomous operational capability. The main contributions of this paper are as follows:
  • We introduce the first hierarchical multi-agent architecture with autonomous task orchestration for pre-consultation, featuring a Controller-based coordination mechanism that enables proactive, structured medical inquiry through intelligent management of specialized agent interactions.
  • We demonstrate superior performance in real-world pre-consultation scenarios, achieving high clinical quality scores across multiple evaluation dimensions on 1,372 clinical cases: Chief Complaint Similarity (4.25), History of Present Illness Similarity (3.71), and Past History Similarity (3.99), significantly outperforming conventional passive inquiry approaches while maintaining workflow coherence.
  • We establish that intelligent coordination mechanisms effectively preserve data privacy and mitigate context limitations in extended dialogues, with our locally deployed framework achieving efficient task completion ($T_{2}$: $12.7$ rounds, $T_{3}$: $16.9$ rounds) and demonstrating 98.2% information collection completeness through adaptive orchestration rather than exhaustive questioning.
Table 1: Statement of Significance
Component Description
Problem Medical consultations globally average under 5 minutes, compromising diagnostic quality. Current AI pre-consultation systems remain passive and struggle with extended dialogue management.
What is Already Known Multi-agent frameworks show promise in healthcare applications, but existing systems lack proactive inquiry capabilities and suffer from context degradation in multi-turn conversations.
What this Paper Adds We introduce the first hierarchical multi-agent architecture with autonomous task orchestration, achieving 98.2% task completion through intelligent coordination and demonstrating superior clinical quality (4.25–4.69/5.0) across 1,372 real cases.
Impact and Implications Healthcare institutions can implement this framework to reduce physician workload while maintaining consultation quality. Researchers gain a model-agnostic architecture for developing proactive medical AI systems. Patients benefit from more thorough pre-consultation assessments.
2 Related Work
2.1 Evolution of Intelligent Agents
The development of intelligent agents based on large language models (LLMs) has undergone rapid evolution from the Transformer architecture [19] replacing recurrence with attention mechanisms, enabling BERT [5] to demonstrate bidirectional context understanding and GPT-3 [3] to achieve emergent few-shot capabilities through large-scale autoregressive pretraining, with subsequent advances in GPT-4 and Gemini [15, 4] showcasing sophisticated reasoning and tool-use abilities. Recent focus has shifted toward collaborative multi-agent systems, with Liu et al. [14] and errag, Tihanyi, and Debbah [6] systematically reviewing LLM-based multi-agent frameworks across domains, while contemporary systems like MegaAgent [22] and AgentNet [25] have introduced decentralized coordination, environmental perception, and long-horizon planning, demonstrating autonomous multi-agent capabilities for complex applications.
2.2 Applications of Intelligent Agents in pre-consultation
Pre-consultation systems have evolved from rule-based dialogue trees to sophisticated multi-agent frameworks, beginning with early work by Bickmore et al.[2] establishing basic conversational patient information collection, Lewis et al.[13] improving adaptability through retrieval-augmented mechanisms integrating patient historical records, Wang and Wong [20] pioneering reinforcement learning-based multi-agent systems with specialized subtask agents, and Tang et al. [18] enhancing clinical fidelity through LLM-powered role-playing agents. Recent advances focus on coordination and contextual management with Prasad et al.[17] introducing on-demand task decomposition, Wang et al.[24] developing priority-based scheduling with dynamic agent generation, and Wang et al.[23] designing ConsultationFlow with modular specialized agents, however existing systems remain fundamentally reactive, lacking proactive inquiry capabilities and struggling with global consistency in complex medical scenarios.
3 Dataset
3.1 Data Collection and Sources
We constructed a comprehensive dataset for medical triage classification by collecting Electronic Health Records (EHRs) from iiyi.com [^1] https://bingli.iiyi.com/, a well-established public medical platform with over 20 years of operational history. This platform serves as a comprehensive EHR sharing repository, encompassing electronic medical records from diverse hospitals across different regions in China. The platform maintains strict privacy standards by applying proper de-identification processes to all shared EHR data, ensuring complete protection of patient privacy information.
Using systematic web crawling techniques, we initially collected 8,172 electronic health records from the platform. To ensure temporal relevance and reflect current medical practices, all collected records were sourced exclusively from the past five years. This temporal constraint guarantees that our dataset captures contemporary medical terminology, treatment approaches, and diagnostic patterns that remain clinically relevant.
3.2 Data Processing and Quality Assurance
We implemented a rigorous two-stage data processing pipeline to ensure high-quality annotations and eliminate inconsistencies. The first stage involved comprehensive quality filtering where we systematically examined all essential fields including History of Present Illness (HPI), Past History (PH), and Chief Complaint (CC). Records were excluded if they contained empty fields, obvious data entry errors, abnormally short HPI descriptions, or cases where HPI and PH content demonstrated excessive similarity that could indicate data duplication or error. This stringent filtering process reduced our initial collection from 8,172 to 2,263 records.
The second stage employed a hybrid validation approach combining automated prediction with expert medical review. We first applied our triage prediction model to each remaining record using the HPI, PH, and CC information as input features. Records where the predicted department assignment matched the original ground truth were automatically validated and retained. For cases where automated prediction disagreed with the original triage decision, we engaged qualified medical professionals to conduct thorough manual verification, ensuring the accuracy and reliability of our ground truth labels.
Our final curated dataset comprises 1,372 validated medical records organized in a hierarchical structure reflecting real-world medical triage systems. The dataset encompasses 9 primary medical departments and 35 secondary sub-specialties, providing comprehensive coverage for both coarse-grained and fine-grained classification tasks. Figure 7 illustrates the distribution of cases across the primary medical departments.
Figure 1: Hierarchical multi-agent framework architecture for medical consultation workflow. Parts 1-3 correspond to Sec. 4.2, 4.3, 4.4 respectively.
4 Agent System
4.1 Framework Overview
In this study, the pre-consultation process $T$ is conceptualized as comprising two stages: Triage and History Taking. The latter, given its large volume of information, heterogeneity of content, and the requirement for structuring doctor–patient communication, can be further organized into three modules: Chief Complaint(CC), History of Present Illness(HPI), and Past History(PH) [7].To achieve systematic information collection and optimization, we formalize the pre-consultation process $T$ as a hierarchical optimization structure comprising four main task groups: Triage ($T_{1}$), HPI collection ($T_{2}$), PH collection ($T_{3}$), and CC generation ($T_{4}$). This hierarchical modeling approach facilitates a balance between global optimality and efficient completion of individual tasks.
$T=$ $\{T_{1},T_{2},T_{3},T_{4}\}$ (1)
$T_{1}=$ $\{t_{11},t_{12}\}$ (2)
$T_{2}=$ $\{t_{21},t_{22},\ldots,t_{26}\}$ (3)
$T_{3}=$ $\{t_{31},t_{32},\ldots,t_{35}\}$ (4)
Based on existing studies[7] and following discussions with multiple physicians, we further decomposed the first three tasks in the pre-consultation process into several domain-specific subtasks (see Table 5). Unlike the decomposable collection tasks, CC generation ($T_{4}$) is a comprehensive integrative task that is not suitable for further subdivision. It operates directly on the Recipient’s incremental recording process, refining information through the continuous accumulation of dialogue content and the dynamic evolution of the HPI. Ultimately, it produces a CC consistent with the full clinical narrative.
Building on the hierarchical task modeling described above, the introduction of a multi-agent framework provides a rational implementation approach. Tasks in the pre-consultation scenario involve both decomposable, structured collection components and context-dependent, integrative generation components, which are often difficult to address simultaneously with a single model. Through the collaborative mechanism of multiple agents, different tasks can be handled by independent autonomous entities while achieving global coordination during interactions, thereby better balancing efficiency and flexibility. The overall architecture is illustrated in Figure 1.
(a) GPT-OSS
(b) Qwen3
(c) Phi4
(d) Uncompleted cases
Figure 2: Performance comparison across different foundation models. Subfigures (a)-(c) show results for GPT-OSS 20B, Qwen3-8B, and Phi4-14B respectively. Subfigure (d) illustrates the distribution of incomplete samples across all three models.
At the implementation level, we designed eight core agents covering global scheduling, question–answer generation, task evaluation, and other processes. This configuration strikes a balance between task coverage and system complexity: the number of agents is sufficient to allow fine-grained task handling while avoiding redundancy and information fragmentation caused by excessive decomposition. Each agent has a clearly defined functional role, forming a structure of role specialization (see Table 6 for the overall design).
To further illustrate the interactions among the eight agents during task execution, we present an overall system architecture diagram (see Figure 1). The diagram is based on the hierarchical task structure and depicts the execution pathways of $T_{1}$, $T_{2}$, $T_{3}$, and $T_{4}$ under multi-agent collaboration. It also visually represents the information flow between global scheduling and local collection processes.
4.2 Task Scheduling and Monitoring
When the system receives the $k$-th round of updated patient information, which includes Triage, History of Present Illness (HPI), Past History (PH), and Chief Complaint (CC), the current task group for evaluation is selected based on the completion order of different sections. If the current task group remains consistent with the previous round’s task group, the system inherits the pending task set $T_{P}$ from the previous round. The Monitor then re-evaluates each subtask within the pending task set $T_{P}$ by incorporating the newly acquired patient information to generate updated Triage, HPI, PH, and CC information. The evaluation is based on two dimensions: clinical semantic validity and task completeness, with scores ranging from 0 to 1. Subtasks that achieve a score above the threshold of 0.85 are removed from the pending task set $T_{P}$, while those scoring below this threshold remain in the set for further processing.
$T_{P}^{k+1}=\{t_{ij}|S(t_{ij})<0.85,t_{ij}\in T_{P}^{k}\}$ (5)
where $S(\cdot)$ denotes the evaluation function of the Monitor, which assesses each subtask based on clinical semantic validity and task completeness.
In this way, the role of the Monitor is not only to evaluate task quality but also to balance information integrity and dialogue efficiency through the threshold mechanism: an excessively high threshold may lead to redundant follow-up questions and over-collection of information, while a threshold that is too low risks missing critical clinical details. Once the Monitor completes its evaluation, the Controller $C$ performs global scheduling based on the real-time state of the pending tasks. It integrates task priority, dialogue context, and the patient’s current clinical information continuously updated by the Recipient. From the pending task set, it selects the subtask with the highest priority as the current subtask $T_{C}$ and generates corresponding inquiry instructions. These instructions are then passed to the Prompter and Inquirer, thereby triggering targeted follow-up questioning.
$T_{C}^{k}=C(T_{p})^{k}$ (6)
4.3 Information Update and Question Generation
Task scheduling alone is insufficient to support a fully interactive consultation. While the scheduling mechanism specifies the current task $T_{C}$, the system must dynamically update patient information during dialogue and generate targeted clinical questions. Without this closed-loop, even accurately assigned tasks cannot ensure the professionality or coherence of the inquiry. To address this, we introduce an information updating and question generation framework atop task scheduling.
Upon receiving natural language expressions from the Virtual Patient, the Recipient integrates the new information into the existing medical record and incrementally updates the Chief Complaint (CC), History of Present Illness (HPI), and Past History (PH). This mechanism allows patient history to evolve across dialogue turns rather than remain at the level of isolated statements, providing a stable context for subsequent tasks. Formally:
$(\mathrm{HPI}_{k+1},\mathrm{PH}_{k+1},\mathrm{CC}_{k+1})=F(\mathrm{HPI}_{k},\mathrm{PH}_{k},\mathrm{CC}_{k},\mathcal{C}_{k+1}),$ (7)
where $\mathcal{C}_{k+1}$ is the complete dialogue up to turn $k+1$, and $F$ denotes the Recipient’s update function. Iterative application ensures progressive accumulation of medical record information.
After updating HPI, PH, and CC, the Prompter generates guiding prompts for the Inquirer to enable more targeted questioning. The Prompter integrates multiple inputs: the updated HPI, PH, and CC from the Recipient, and importantly, the current task $T_{C}$ provided by the Controller. By incorporating $T_{C}$, the Prompter ensures that the generated prompts align precisely with the current task objectives while remaining within the scope of clinician–patient interactions, avoiding irrelevant or redundant queries. In effect, the Prompter operationalizes the abstract goals of task scheduling.
Guided by the prompts from the Prompter and the updated medical record, the Inquirer produces context-sensitive clinical questions. For a given current task $T_{C}$, the Inquirer’s outputs reflect the scheduling objectives while incorporating the constructed clinical context, maintaining both logical continuity and medical professionalism.
Additionally, when task $T_{1}$ has not been completed, the Triager is invoked as an auxiliary component to analyze symptom characteristics and provide triage recommendations. It constructs a contextualized clinical interpretation based on structured taxonomies and outputs candidate recommendations for primary and secondary departments. This maps unstructured dialogue into a clinically interpretable feature space during the initial triage phase. Once $T_{1}$ is completed, the Triager is no longer called in subsequent rounds.
4.4 Evaluation
Within the entire multi-agent framework, the Evaluator is designed as an independent assessment agent responsible for systematically measuring system performance. It is important to emphasize that the Evaluator does not participate in the actual clinician–patient dialogue and does not influence task scheduling or subtask selection; its function is entirely limited to post hoc evaluation. Thus, it serves a role of performance characterization and quality feedback rather than interactive execution. To ensure comprehensive and rigorous evaluation, the Evaluator operates on two complementary levels: dialogue interaction and content accuracy. This distinction reflects the dual objectives of clinical pre-consultation: ensuring both the rationality of the interaction process and the accuracy of the generated content.
At the interaction level, it examines whether the system’s questions follow clinical logic and task objectives, the clarity and professionalism of language, the completeness of information collection, and the overall professionalism demonstrated during the inquiry. At the content level, it evaluates the alignment of the three key medical record elements—Chief Complaint, History of Present Illness, and Past History—against reference standards to ensure consistency and fidelity. Detailed definitions and criteria for each dimension are provided in Table 7.
Regarding scoring methodology, the Evaluator uses a 0–5 scale for all seven dimensions, where 0 represents severe omission or error, and 5 represents an ideal clinical level. Complementarily, qualitative criteria are provided to interpret each score range with specific explanations and examples. For instance, Table 8 presents the scoring standards for Past History Similarity (PHS), considering information coverage, chronological consistency, and accuracy of medical terminology. By combining quantitative and qualitative assessments, the Evaluator not only ensures objective scoring but also provides interpretable feedback for subsequent system optimization.
5 Experiment
5.1 Triage Experiment
To evaluate the effectiveness of our multi-agent framework in medical triage scenarios, we conducted comprehensive experiments on department classification tasks. We assessed the system’s ability to correctly route patients to both primary and secondary departments based on their presenting symptoms and medical history.
Table 2 presents the classification accuracy across multiple iteration steps. Our results demonstrate consistent improvement in triage accuracy through iterative refinement. The primary department classification accuracy increased from 83.0% in the initial step to 87.0% by step 4, representing a 4.0 percentage point improvement. Similarly, secondary department classification showed substantial gains, rising from 75.4% to 80.5%, an improvement of 5.1 percentage points. The progressive improvement across iterations validates our multi-agent approach, where agents collaboratively refine their decisions through structured dialogue and knowledge sharing.
Table 2: Performance progression across steps for primary and secondary department classification. Results show accuracy improvements through multi-round iteration.
Iteration Step
Primary
Department
Secondary
Department
Step 1 83.0% 75.4%
Step 2 85.6% 78.7%
Step 3 86.4% 80.2%
Step 4 87.0% 80.5%
Figure 3 illustrates the triage accuracy distribution across different medical departments. We observe significant variability in performance across specialties, with accuracy ranging from approximately 65% to 95%. Ophthalmology achieved the highest classification accuracy at 94.8%, likely due to the distinctive nature of ocular symptoms that facilitate clear differentiation. In contrast, Psychiatry exhibited the lowest accuracy at 65.2%. Further analysis revealed that misclassified psychiatry cases were predominantly routed to the neurology department. This pattern likely reflects the inherent overlap between neurological and psychiatric presentations, as well as potential variations in departmental structures across different hospital systems.
Figure 3: Triage accuracy across different medical departments. The bar chart shows significant variation in classification performance, with Ophthalmology achieving the highest accuracy (100.0%) and Psychiatry the lowest (30.0%).
These findings demonstrate the value of our multi-agent architecture in handling the inherent ambiguity of medical triage. The consistent improvement through iterative refinement and the ability to achieve high accuracy for departments with distinct symptom profiles suggest our system could provide immediate value in clinical settings, particularly in routing clear-cut cases and flagging ambiguous cases for human review.
5.2 Experimental Design
All experiments in this study were conducted based on our self-constructed dataset comprising 1,372 validated medical records. These records serve as the foundation for evaluating the performance of our proposed consultation system across various clinical scenarios. For model deployment, all models were deployed on a single NVIDIA A100 GPU using GPUStack as the deployment framework. To enhance inference efficiency, we implemented concurrent processing with 60 parallel threads, allowing the system to handle multiple consultation sessions simultaneously while maintaining response quality. The maximum number of dialogue turns was set to 30 rounds per consultation session. This constraint ensures practical applicability while providing sufficient opportunity for comprehensive information gathering. Samples that failed to complete all three task groups ($T_{1}$, $T_{2}$, and $T_{3}$) within the 30-round limit were classified as failures in our evaluation metrics. To ensure consistency and fairness in evaluation across all experiments, we standardized the Evaluator component by uniformly employing the GPT-OSS 20B model. This standardization eliminates potential variations that could arise from using different evaluation models, thereby ensuring reliable and comparable results throughout our experimental analysis.
5.3 Task completion across different models
To evaluate the generalizability of the hierarchical multi-agent architecture, we conducted zero-shot tests on three LLMs of different scales: GPT-OSS 20B, Qwen3-8B, and Phi4-14B. The results (Figure 2) show that our framework can guide all models to complete multi-turn medical consultations without any task-specific fine-tuning, highlighting its model-agnostic nature.
Specifically, the powerful GPT-OSS 20B achieved the most stable and reliable performance under the framework’s coordination, with the lowest number of unfinished cases. The combined unfinished count for $T_{2}$ and $T_{3}$ tasks by GPT-OSS 20B was only 55.43% of that achieved by Qwen3-8B and merely 8.08% of Phi4-14B’s unfinished cases, demonstrating superior task completion capability. Interestingly, Qwen3-8B demonstrated remarkable efficiency in task execution. It completed both $T_{2}$ and $T_{3}$ tasks with fewer average rounds compared to GPT-OSS 20B. This efficiency advantage may be attributed to Qwen3-8B’s stronger Chinese language capabilities, which are particularly relevant for our medical consultation scenarios. However, its lower completion rate compared to GPT-OSS 20B suggests that superior reasoning abilities remain crucial for ensuring task success. For Phi4-14B, the overall performance lagged behind both GPT-OSS 20B and Qwen3-8B, which may be related to its earlier release date (Phi4-14B was released in 2024, while GPT-OSS 20B and Qwen3-8B were released in 2025). Despite this limitation, the framework still exhibited strong robustness and fault tolerance, successfully guiding Phi4-14B to complete a considerable number of cases rather than failing entirely.
All models required more rounds to complete the more complex $T_{3}$ tasks, which is consistent with expectations and further validates the soundness of the evaluation design. The experimental results demonstrate that the system maintains effectiveness across different models, and notably, task completion success rate does not solely depend on model parameter size but rather correlates more strongly with language-specific capabilities and general reasoning abilities.
5.4 Evaluation scores across different models
As illustrated, all three models achieved commendable evaluation scores across the six distinct medical consultation tasks (CI, CQ, IC, OP, PHS, CCS). The consistent presence of high median values (black line within each box) and the tight interquartile ranges (compact box size) indicate that our architecture effectively guides different LLMs to perform the structured reasoning required for multi-round dialogues.
In the three metrics that reflect information extraction effectiveness—Chief Complaint Similarity (CCS), Present Illness Similarity (HPIS), and Past History Similarity (PHS)—all three models (GPT-OSS 20B, Qwen3-8B, and Phi4-14B) achieved average scores of at least 3.74, as shown in Figure 4. Specifically, GPT-OSS 20B achieved an average of 3.99, Qwen3-8B scored 3.74, and Phi4-14B also reached 3.99. These consistently high scores demonstrate that our system can achieve excellent final outcomes across different models, regardless of their individual characteristics.
Regarding the four metrics that reflect dialogue process quality (CI, CQ, IC, and OP), Phi4-14B demonstrated slightly stronger performance compared to GPT-OSS 20B and Qwen3-8B. This performance variation may be attributed to differences in the linguistic styles of different models, suggesting that certain conversational characteristics inherent to Phi4-14B align well with the dialogue evaluation criteria.
Figure 4: Mean scores and variance across seven evaluation dimensions for GPT-OSS 20B, Qwen3-8B, and Phi4-14B. Error bars indicate standard deviation.
5.5 Completion rate across different scheduling strategies
We compare the dynamic performance of two representative scheduling strategies: Default Order (baseline sequential prompting following a predefined sequence, which we designed with reference to clinical importance) and Agent Driven (autonomous coordination through agent selection).
As shown in Fig. 5, both strategies exhibit a monotonic upward trend in task completion across the 13 subtasks. However, they differ in convergence levels and growth rates.
Specifically, the Agent Driven strategy demonstrates the fastest growth in the early phase and ultimately stabilizes at the highest completion rate of 98.2% around round 25, reflecting superior global coordination and convergence capacity. In contrast, the Default Order strategy maintains a steadier growth trajectory during the middle phase (rounds 5–15) but plateaus at a lower final completion rate of 93.1%, suggesting a performance ceiling.
These results indicate that the Agent Driven strategy achieves a more favorable balance between stability and efficiency, enabling faster and more complete subtask fulfillment. Meanwhile, the Default Order approach provides moderate consistency but lacks the capacity to overcome performance saturation. The superior performance of Agent Driven strategy demonstrates the advantage of dynamic, intelligent scheduling over static sequential processing, even when the default sequence is designed with clinical considerations in mind.
(a) Agent Driven
(b) Medical Priority
Figure 5: Task completion rates under different scheduling strategies. Our proposed Agent Driven strategy demonstrates superior completion performance compared to the baseline Medical Priority approach.
5.6 Evaluation scores across different scheduling strategies
To further assess the impact of scheduling strategies on medical history generation, we compared Default Order and Agent Driven across six clinically relevant dimensions: Chief Complaint Similarity (CCS), Present Illness Similarity (HPIS), Past History Similarity (PHS), Inquiry Consistency (IC), Overall Progress (OP), and Clinical Integrity (CI).
As illustrated in Fig. 6, the Agent Driven strategy consistently outperforms Default Order across most dimensions. It achieves notable gains in IC and OP, reflecting stronger coherence and broader coverage, and shows clear advantages in PHS and HPIS, suggesting more complete and accurate medical history collection. While the two strategies achieve comparable performance in CCS, Agent Driven still maintains a slight edge. The only area where Default Order remains competitive is CI, though its overall balance lags behind.
Taken together, these results highlight the robustness of Agent Driven scheduling: it not only accelerates task completion (as shown in the previous subsection) but also produces more consistent and comprehensive outputs across diverse evaluation metrics. This demonstrates that dynamic, intelligent coordination surpasses static sequential processing, even when the default sequence is carefully designed based on clinical considerations.
Figure 6: Evaluation scores across different scheduling strategies
5.7 Real-World Evaluation
Although the Evaluator-based scoring offers consistency and reference value, its evaluation focus may differ from real-world pre-consultation settings. To validate the usability and clinical relevance of the results generated by the Recipient, we incorporate expert physician evaluation as a complementary metric. A total of $40$ samples were randomly selected from a dataset of $1,372$ records. For each sample, considering that lengthy dialogues may cause evaluation fatigue and affect scoring consistency, we randomly sampled $5$ out of the $20$ dialogue turns to present to evaluators. This sampling strategy balances the need for comprehensive evaluation while maintaining physician focus and reducing cognitive load. The sampled dialogue turns, along with the final-round generated Chief Complaint (CC), History of Present Illness (HPI), and Past History (PH) were presented. A total of $18$ licensed physicians were invited to participate in the evaluation, with each physician independently assessing $20$ samples. Each sample was thus evaluated by $9$ physicians, who assigned a score on a discrete scale from $1$ to $5$.
Table 3: Real World Evaluation
Field Average Score Standard Deviation
$C$ 4.5647 0.2221
HPI 4.4794 0.1284
PH 4.6853 0.0997
Table 3 presents the expert physician evaluations on generated content across different fields. The average scores are above $4.45$, indicating generally high quality, while relatively low standard deviations (below $0.23$) demonstrate consistent agreement among raters.
Acknowledgments
The authors gratefully acknowledge the financial support provided by the Guangdong Provincial Department of Science and Technology through Grant 2023CX10X070, the Guangdong Provincial Key Laboratory of Human Digital Twin through Grant 2022B1212010004, the Guangzhou Basic Research Program through Grant SL2023A04J00930, and the Shenzhen Holdfound Foundation Endowed Professorship.
Appendix A Case Study and Human Evaluation
Next, we present a representative case study featuring a 5-round dialogue transcript along with the corresponding generated HPI, PH, and CC outputs. Additionally, we provide evaluation scores from nine medical practitioners for this case to better demonstrate the practical effectiveness of our proposed model.
A.1 Dialogue Excerpt
Patient-Doctor Consultation Transcript Patient: My neck and shoulder have been hurting for half a month. It’s hard to move. Doctor: Please think back carefully. When the pain started, did anything specific happen? Patient: That night, I got up too quickly while turning over, and suddenly my right neck cramped with sharp pain. After resting for 10 minutes, it eased. Doctor: Can you describe the nature of the pain? Is it a dull constant ache, intermittent stabbing, or something else? Patient: It’s mostly spasmodic pain in the neck and right shoulder. Very intense during attacks, but subsides after rest. The left shoulder feels sore and stiff in the morning. Doctor: Any other symptoms, such as numbness in the arm, dizziness, or vision changes? Patient: No, just the neck and shoulder pain. Doctor: Have you ever been diagnosed with any chronic diseases, like hypertension, diabetes, heart disease, hepatitis, etc.? Patient: No, I’ve always been healthy. Doctor: Any allergies to drugs or foods? Patient: No known allergies.
A.2 Generated Documentation
System-Generated Medical Documentation Chief Complaint ($C$):
Neck and shoulder pain with limited mobility for half a month.
History of Present Illness (HPI):
Patient developed neck and shoulder pain with restricted mobility half a month ago. One day ago, a sudden movement while getting up led to acute cramping pain in the right neck, which resolved after 10 minutes of rest. The pain is intense during attacks and alleviates with rest. Left shoulder shows noticeable morning soreness and limited mobility. Denies numbness or dizziness.
Past History (PH):
Denies any history of chronic diseases such as hypertension or diabetes. No prior surgeries. No known drug or food allergies.
A.3 Human Evaluation Results
Nine medical practitioners evaluated the generated documentation quality, with scores shown in Table 4.
Table 4: Expert evaluation scores on one representative case
Field 1 pt 2 pt 3 pt 4 pt 5 pt Avg. Score
$C$ 0 0 0 2 7 4.78
HPI 0 0 0 3 6 4.67
PH 0 0 0 0 9 5.00
Appendix B Supplementary Tables and Figures
This appendix provides detailed task definitions and additional visualizations that complement the main text. The hierarchical task decomposition framework (Table 5) presents the complete structure of our pre-consultation system, while supplementary figures and tables offer comprehensive evaluation criteria and data distributions.
B.1 Task Decomposition Framework
The following table presents the complete hierarchical decomposition of pre-consultation tasks, which forms the foundation of our multi-agent coordination system.
Table 5: Task definitions in the pre-consultation framework
Task Subtasks Description
Triage Primary Department Identification Determine the primary department the patient should visit.
Secondary Department Identification Identify the specific secondary department based on the primary department.
[0.5pt/2pt] HPI Collection Onset Record the time, location, mode of onset, prodromal symptoms, and possible causes or triggers.
Main Symptom Characteristics Describe the location, nature, duration, severity, and aggravating/relieving factors of main symptoms in chronological order.
Disease Progression Describe the progression and evolution of the illness in chronological order.
Accompanying Symptoms Record accompanying symptoms and their relationship with the main symptoms.
Diagnostic and Therapeutic History Record whether the patient has undergone examinations or treatments after onset, and their outcomes if applicable.
General Condition Briefly record the patient’s mental state, sleep, appetite, bowel and bladder functions, and body weight after onset.
[0.5pt/2pt] PH Collection Disease History Record the patient’s history of past illnesses, including infectious diseases such as tuberculosis and hepatitis.
Immunization History Inquire about the patient’s vaccination history.
Surgical and Trauma History Record the patient’s history of surgeries and traumas.
Blood Transfusion History Inquire about the patient’s history of blood transfusions and any adverse reactions.
Allergy History Inquire about the patient’s history of food or drug allergies.
B.2 Data Distribution and System Architecture
Figure 7 illustrates the distribution of medical cases across primary departments in our validation dataset, demonstrating the comprehensive coverage of clinical specialties.
Figure 7: Distribution of clinical departments in our dataset. The chart shows the proportional representation of primary medical specialties, with Internal Medicine comprising the largest share (38.4%) and Ophthalmology the smallest (0.3%).
B.3 Agent Architecture and Evaluation Framework
The following tables provide detailed specifications of our multi-agent system components and evaluation criteria used throughout the experimental validation.
Table 6: Overview of the Agents and Their Functions
Agent Name Function
Virtual Patient Generate realistic clinical presentations based on validated medical records
Recipient Update and maintain evolving medical records throughout dialogue sessions
Triager Perform hierarchical department-level triage with primary and secondary classification
Monitor Assess subtask completion status using clinical validity and completeness metrics
Controller Select optimal next subtask through priority-based scheduling algorithm
Prompter Formulate context-aware inquiry strategies based on current task objectives
Inquirer Produce clinically appropriate medical questions maintaining dialogue coherence
Evaluator Provide comprehensive performance assessment across multiple quality dimensions
Table 7: Evaluation Dimensions for the Evaluator
Dimension Description
Clinical Inquiry (CI) Assess completeness, professionalism, and logical flow of medical history collection process
Communication Quality (CQ) Evaluate fluency, clarity, and empathy in doctor-patient dialogue interactions
Information Completeness (IC) Measure the integrity, systematicity, and focus of information collection across all domains
Overall Professionalism (OP) Assess domain knowledge accuracy, appropriate terminology usage, and structured clinical reasoning
CC Similarity (CCS) Compare generated chief complaint content and semantics with reference standard
HPI Similarity (HPIS) Measure coverage and structural alignment compared to reference History of Present Illness
PH Similarity (PHS) Evaluate correspondence with reference Past History including completeness and accuracy
Table 8: Scoring Criteria for Past History Similarity Assessment
Score Description
0 Past History not mentioned or contains insufficient information for meaningful assessment
1 Partially accurate; loosely related to true PH but contains notable omissions or inaccuracies
3 Generally accurate; covers most key elements of true PH with minor deviations or missing details
5 Highly accurate; closely matches true PH with comprehensive coverage and precise medical terminology